Please explain the concept of surface tension. How does it act like a rubbered sheet? If it does, then why does it tend to occupy less surface area? Please explain in the most simplified explanation.

Grade:

1 Answers

Abhijeet Khalkar
13 years ago

Surface tension is a property of the surface of a liquid that allows it to resist an external force. It is revealed, for example, in floating of some objects on the surface of water, even though they are denser than water, and in the ability of some insects and even reptiles to run on the water surface. This property is caused by cohesion of like molecules, and is responsible for many of the behaviors of liquids.

Surface tension has the dimension of force per unit length, or of energy per unit area. The two are equivalent—but when referring to energy per unit of area, people use the term surface energy—which is a more general term in the sense that it applies also to solids and not just liquids.

In materials science, surface tension is used for either surface stress or surface free energy.

 

This equation is known as the Young–Laplace equation

\Delta p\ =\ \gamma \left( \frac{1}{R_x} + \frac{1}{R_y} \right)

                 where:

  • Δp is the pressure difference.
  • \scriptstyle\gamma is surface tension.
  • Rx and Ry are radii of curvature in each of the axes that are parallel to the surface.

 

Surface tension can be measured using the pendant drop method on a goniometer.